OpenStudy (decarr432):

3a^3-5b^3+?a^3+?b^3=(a^3+b^3

OpenStudy (decarr432):

\[(3a^3-5b^3)+?a^3+?b^3=(a^3+b^3\]

OpenStudy (decarr432):

I think the answer it -3a^3 5b^3

zepdrix (zepdrix):

What are the coefficients on the right? :) \(\large\rm a^3=?~~~0a^3\) What is the magical number that they're hiding from us? Is it a zero?

OpenStudy (decarr432):

I have to find two numbers one for a and one for b that make the equation true to equal (a^3b^3)

OpenStudy (decarr432):

+

OpenStudy (alivejeremy):

thats the answer

zepdrix (zepdrix):

I'm just trying to get you to realize how many a^3 are on the right side of the equation.

OpenStudy (decarr432):

yeah 0 for a and b

zepdrix (zepdrix):

No no no. 0apples is just 0, I wouldn't even have to write the apples. When I write the word "apple", that implies ONE apple, yes? When I don't write a number next to it, it's an invisible 1, not an invisible 0 :)

zepdrix (zepdrix):

So you have 1a^3 and 1b^3 on the right side of your equation.

OpenStudy (decarr432):

okay I see now the answer you be a=-2 b=+6

OpenStudy (decarr432):

was just one off

zepdrix (zepdrix):

Ahh there we go! :)

OpenStudy (decarr432):

This is why you are the best!

zepdrix (zepdrix):

lol :p